### 07 Mar 2024 v1

# On semidefinite descriptions for convex hulls of quadratic programs

Abstract
Quadratically constrained quadratic programs (QCQPs) are a highly expressive class of nonconvex optimization problems. While QCQPs are NP-hard in general, they admit a natural convex relaxation via the standard semidefinite program (SDP) relaxation. In this paper we study when the convex hull of the epigraph of a QCQP coincides with the projected epigraph of the SDP relaxation. We present a sufficient condition for convex hull exactness and show that this condition is further necessary under an additional geometric assumption. The sufficient condition is based on geometric properties of $\Gamma$, the cone of convex Lagrange multipliers, and its relatives $\Gamma_{1}$ and $\Gamma^{\circ}$.
1 Introduction
A quadratically constrained quadratic program (QCQP) is an optimization problem of the form
$\operatorname{Opt}\coloneqq\inf_{x\in{\mathbb{R}}^{n}}\left\{q_{\textup{obj}}(x):\,q_{i}(x)\leq 0,\,\forall i\in[m]\right\},$ (1)
where $q_{\textup{obj}},q_{1},\dots,q_{m}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ are each (possibly nonconvex) quadratic functions. For each $i\in\left\{{\textup{obj}}\right\}\cup[m]$, we will write $q_{i}(x)=x^{\top}A_{i}x+2b_{i}^{\top}x+c_{i}$ for $A_{i}\in{\mathbb{S}}^{n}$, $b_{i}\in{\mathbb{R}}^{n}$, and $c_{i}\in{\mathbb{R}}$. These optimization problems arise naturally in a variety of application areas (see e.g., [25] and references therein).
Although QCQPs are NP-hard in general, they admit a natural tractable convex relaxation known as the standard (Shor) semidefinite program (SDP) relaxation,
$\operatorname{Opt}_{\textup{SDP}}\coloneqq\inf_{x\in{\mathbb{R}}^{n}}\left\{\left\langle A_{\textup{obj}},X\right\rangle+2b_{\textup{obj}}^{\top}x+c_{\textup{obj}}:\,\begin{array}[]{l}\exists X\succeq xx^{\top}:\\ \left\langle A_{i},X\right\rangle+2b_{i}^{\top}x+c_{i}\leq 0,\,\forall i\in[m]\end{array}\right\}.$
A recent and growing line of work has offered deterministic conditions under which the SDP relaxation of a general QCQP is exact for various definitions of exactness. In their celebrated paper, Fradkov and Yakubovich [8] prove the S-lemma, which implies that the problem of minimizing an arbitrary quadratic objective function over the unit ball (or any single quadratic constraint) can be solved via SDP techniques. Specifically, the S-lemma implies that objective value exactness—the condition that the optimal value of the QCQP and the optimal value of its SDP relaxation coincide—holds for QCQPs with a single constraint. Locatelli [16], Jeyakumar and Li [14], Bomze et al. [4] present sufficient conditions for objective value exactness to continue to hold in the presence of additional linear constraints and/or one additional quadratic constraint (possibly using strengthened relaxations). Burer and Ye [6] study diagonal QCQPs—those QCQPs for which $A_{\textup{obj}},A_{1},\dots,A_{m}$ are diagonal matrices—with a general number of constraints and give sufficient conditions for objective value exactness. Sojoudi and Lavaei [21] study sparse SDPs where the $A_{i}$ matrices have an underying graph structure and give sufficient conditions for objective value exactness phrased in terms of the graph structure and the associated coefficients. Wang and Kılınç-Karzan [25] continue this line of work by developing a general framework for deriving sufficient conditions for both objective value exactness and convex hull exactness—the condition that the convex hull of the QCQP epigraph coincides with the (projected) SDP epigraph—for QCQPs where a specific dual set, $\Gamma_{1}$, is polyhedral (see Section 2). Beyond being a natural sufficient condition for objective value exactness, convex hull exactness has its own applications and motivation, particularly in the derivation of strong relaxations of certain critical substructures in nonconvex problems; see e.g., [23, 25] and references therein.
While the framework presented in [25] at once covers and extends many existing results on exactness [15, 6, 8, 12, 18, 26, 5], it is still quite limited. In particular, the assumption that $\Gamma_{1}$ is polyhedral is rarely satisfied outside of simultaneously diagonalizable QCQPs and precludes the results in [25] from being applicable to a wider range of interesting QCQPs. In this paper, we generalize the results introduced in [25] by eliminating its reliance on the polyhedrality assumption. Specifically, after formally defining our setup and recalling basics on the SDP relaxation of a QCQP, and the cone of convex Lagrange multipliers $\Gamma$ (and the closely related sets $\Gamma_{1}$ and $\Gamma^{\circ}$) in Section 2, we present in Section 3 a sufficient condition for convex hull exactness that generalizes [25, Theorem 1]. This sufficient condition (Theorem 1) states that convex hull exactness holds as long as certain systems of equations (that depend on $\Gamma$, $\Gamma_{1}$, or $\Gamma^{\circ}$) contain nontrivial solutions. Thus, this sufficient condition can be checked in a systematic manner whenever $\Gamma$, $\Gamma_{1}$, or $\Gamma^{\circ}$ is sufficiently simple. Furthermore, in Theorem 2, we show that our sufficient condition for convex hull exactness is also necessary under the assumption that $\Gamma^{\circ}$ is facially exposed (see Assumption 2 and its surrounding discussion). To the best of our knowledge, this is the first necessary and sufficient condition for convex hull exactness even in the context of diagonal QCQPs (where $\Gamma$ is automatically polyhedral). Based on these results, we also revisit the previous polyhedral setting of [25] and provide more refined necessary and sufficient conditions for convex hull exactness in this setting (see Sections 3.4 and 5). These results offer an algebraic way to verify convex hull exactness even when the QCQP is not simultaneously diagonalizable. We illustrate our results (Theorems 1 and 3.4) on three applications (a prototypical mixed binary set arising in sparse regression problems, quadratic matrix programs (QMPs), and the sets defined by two quadratic constraints) in Appendix B. In the first two applications, the resulting $\Gamma$ sets are non-polyhedral, and thus the sufficient conditions from [25] that work under the polyhedrality assumption of $\Gamma$ are not applicable. For QMPs, we show that convex hull exactness holds whenever the number of constraints is small (when compared to the rank of the matrix variable) which strengthens the objective value exactness result of [2]. In the last application, we illustrate that our Section 3.4 reduces to a succinct and easy to check necessary and sufficient condition for convex hull exactness when $q_{\textup{obj}}=0$ and $m=2$ as opposed to the more difficult to check condition of [26].
1.1 Notation
For a positive integer $n$, let $[n]\coloneqq\left\{1,\dots,n\right\}$. For $i\in[n]$, let $e_{i}$ be the $i$th standard basis vector in ${\mathbb{R}}^{n}$. Let $0_{n}$ denote the zero vector in ${\mathbb{R}}^{n}$. Let ${\mathbb{S}}^{n}$ denote the vector space of $n\times n$ real symmetric matrices and ${\mathbb{S}}^{n}_{+}$ the cone of positive semidefinite matrices. For $M\in{\mathbb{S}}^{n}$, we write $M\succeq 0$ (resp. $M\succ 0$) to denote that $M$ is positive semidefinite (resp. positive definite). Let $\ker(M)$ denote the kernel of $M$. Suppose ${\cal M}$ is a subset of some Euclidean space. Let $\operatorname{conv}({\cal M})$, $\operatorname{clcone}({\cal M})$, $\operatorname{span}({\cal M})$, ${\cal M}^{\perp}$, $\operatorname{rint}({\cal M})$, and $\dim({\cal M})$ denote the convex hull, closed conic hull, span (linear hull), orthogonal complement, relative interior, and dimension of ${\cal M}$ respectively. Specifically, if $\operatorname*{\mathbb{E}}$ is the Euclidean space containing ${\cal M}$, then ${\cal M}^{\perp}\coloneqq\left\{x\in\operatorname*{\mathbb{E}}:\,\left\langle x,m\right\rangle=0,\,\forall m\in{\cal M}\right\}$. Suppose $K$ is a cone in some Euclidean space. Let $K^{\circ}$ denote the polar cone of $K$. The notation $F\trianglelefteq K$ denotes that $F$ is a face of $K$. By convention, faces of cones are always nonempty. Given $x,y$ in some Euclidean space $\operatorname*{\mathbb{E}}$, let $[x\pm y]\coloneqq\operatorname{conv}(\left\{x-y,x+y\right\})$.
2 Preliminaries
We study the epigraph of the QCQP in (1) defined as
${\cal S}\coloneqq\left\{(x,t)\in{\mathbb{R}}^{n}\times{\mathbb{R}}:\,\begin{array}[]{l}q_{\textup{obj}}(x)\leq 2t\\ q_{i}(x)\leq 0,\,\forall i\in[m]\end{array}\right\}.$
It is well known in the QCQP literature [3, 25, 10] that the SDP relaxation of a QCQP is equivalent (under standard assumptions) to the dual of the Lagrangian dual. We start by introducing notation related to Lagrangian aggregation and then state this equivalence formally in Remark 1.
Let $q:{\mathbb{R}}^{n}\to{\mathbb{R}}^{1+m}$ be indexed by $\left\{{\textup{obj}}\right\}\cup[m]$ where $q(x)_{\textup{obj}}=q_{\textup{obj}}(x)$ and $q(x)_{i}=q_{i}(x)$ for $i\in[m]$. Let $e_{\textup{obj}},e_{1},\dots,e_{m}$ denote the corresponding unit vectors in ${\mathbb{R}}^{1+m}$. Note that for $(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}^{1+m}$, the expression $\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle=\gamma_{\textup{obj}}q_{\textup{obj}}(x)+\sum_{i=1}^{m}\gamma_{i}q_{i}(x)$ is a quadratic function in $x\in{\mathbb{R}}^{n}$. Define $A(\gamma_{\textup{obj}},\gamma)\coloneqq\gamma_{\textup{obj}}A_{\textup{obj}}+\sum_{i\in[m]}\gamma_{i}A_{i}$. Similarly, define $b(\gamma_{\textup{obj}},\gamma)$, and $c(\gamma_{\textup{obj}},\gamma)$.
We will at times work on the slice of ${\mathbb{R}}^{1+m}$ where the coordinate $\gamma_{\textup{obj}}$ is fixed to $\gamma_{\textup{obj}}=1$. Let $A[\gamma]\coloneqq A(1,\gamma)$ and similarly define $b[\gamma]$ and $c[\gamma]$. Set $[\gamma,q(x)]\coloneqq\left\langle(1,\gamma),q(x)\right\rangle=q_{\textup{obj}}(x)+\sum_{i=1}^{m}\gamma_{i}q_{i}(x)$.
Note that
$\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle=x^{\top}A(\gamma_{\textup{obj}},\gamma)x+2b(\gamma_{\textup{obj}},\gamma)^{\top}x+c(\gamma_{\textup{obj}},\gamma),\quad\text{and}$
$[\gamma,q(x)]=x^{\top}A[\gamma]x+2b[\gamma]^{\top}x+c[\gamma].$
We recall and extend the following definition from [25].
Definition 1.
We define $\Gamma$ as the cone of convex Lagrange multipliers for (1), i.e.,
$\Gamma\coloneqq\left\{(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}_{+}\times{\mathbb{R}}^{m}_{+}:\,A(\gamma_{\textup{obj}},\gamma)\succeq 0\right\},\qquad\text{and also its cross section}$
$\Gamma_{1}\coloneqq\left\{\gamma\in{\mathbb{R}}_{+}^{m}:\,(1,\gamma)\in\Gamma\right\}=\left\{\gamma\in{\mathbb{R}}_{+}^{m}:\,A[\gamma]\succeq 0\right\}.$
These sets are related to the feasible domain of a partial dual of the SDP relaxation (see Remark 2). Note that given $(\gamma_{\textup{obj}},\gamma)\in\Gamma$, the quadratic function $x\mapsto\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle$ is convex. Similarly, given $\gamma\in\Gamma_{1}$, the quadratic function $x\mapsto[\gamma,q(x)]$ is convex.
We will make the following blanket assumption for the remainder of the paper. This assumption can be interpreted as a dual strict feasibility condition and is standard in the literature [6, 25].
Assumption 1.
There exists $(\gamma_{\textup{obj}}^{*},\gamma^{*})\in\Gamma$ such that $A(\gamma_{\textup{obj}}^{*},\gamma^{*})\succ 0$.
Assumption 1 is equivalent to the assumption that there exists $\gamma^{*}\in\Gamma_{1}$ such that $A[\gamma^{*}]\succ 0$. One direction is immediate. In the other direction, suppose $A(\gamma^{*}_{\textup{obj}},\gamma^{*})\succ 0$. As the minimum eigenvalue of $A(\gamma^{*}_{\textup{obj}},\gamma^{*})$ is continuous in its arguments, we may assume that $\gamma^{*}_{\textup{obj}}>0$. Then, $A[\gamma^{*}/\gamma^{*}_{\textup{obj}}]=\frac{1}{\gamma^{*}_{\textup{obj}}}A(\gamma^{*}_{\textup{obj}},\gamma^{*})\succ 0$.
Remark 1.
Under Assumption 1, $\Gamma$ is the closed cone generated by its slice at $\gamma_{\textup{obj}}=1$, i.e., $\Gamma=\operatorname{clcone}\left(\left\{(1,\gamma):\,\gamma\in\Gamma_{1}\right\}\right)$. (See discussion following [25, Assumption 2].)
The projected SDP relaxation of ${\cal S}$ is given by
${\cal S}_{\textup{SDP}}\coloneqq\left\{(x,t)\in{\mathbb{R}}^{n}\times{\mathbb{R}}:\,\begin{array}[]{l}\exists X\succeq xx^{\top}:\\ \left\langle A_{\textup{obj}},X\right\rangle+2b_{\textup{obj}}^{\top}x+c_{\textup{obj}}\leq 2t\\ \left\langle A_{i},X\right\rangle+2b_{i}^{\top}x+c_{i}\leq 0,\,\forall i\in[m]\end{array}\right\},$ (5)
and $\operatorname{Opt}_{\textup{SDP}}\coloneqq\inf_{(x,t)\in{\cal S}_{\textup{SDP}}}2t$. By taking $X=xx^{\top}$ in (5), we see that $\operatorname{Opt}\geq\operatorname{Opt}_{\textup{SDP}}$ and $\operatorname{conv}({\cal S})\subseteq{\cal S}_{\textup{SDP}}$.
The following lemma rewrites ${\cal S}_{\textup{SDP}}$ in terms of $\Gamma$, $\Gamma_{1}$, and $\Gamma^{\circ}$ and follows from a straightforward duality argument; see Figure 1 for an illustration of these sets. These expressions allow for a straightforward analysis of ${\cal S}_{\textup{SDP}}$ whenever the corresponding dual set is sufficiently simple. {restatable}lemmalemSSDP Suppose Assumption 1 holds. Then,
${\cal S}_{\textup{SDP}}$ =$=\left\{(x,t)\in{\mathbb{R}}^{n+1}:\,[\gamma,q(x)]\leq 2t,\,\forall\gamma\in\Gamma_{1}\right\}$
=$=\left\{(x,t)\in{\mathbb{R}}^{n+1}:\begin{array}[]{r}\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle\leq 2\gamma_{\textup{obj}}t,\,\forall(\gamma_{\textup{obj}},\gamma)\in\Gamma\end{array}\right\}$
=$=\left\{(x,t)\in{\mathbb{R}}^{n+1}:\,q(x)-2te_{\textup{obj}}\in\Gamma^{\circ}\right\}.$
The proof of Remark 1 is given in Appendix A. We list some corollaries of Remark 1.
Corollary 1.
Suppose Assumption 1 holds. Then, $\operatorname{Opt}_{\textup{SDP}}=\inf_{x\in{\mathbb{R}}^{n}}\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$.
Corollary 2.
Suppose Assumption 1 holds. Then, ${\cal S}_{\textup{SDP}}$ is closed.
Remark 2.
One may compute $\operatorname{Opt}_{\textup{SDP}}$ in the form (5) by minimizing over $x\in{\mathbb{R}}^{n}$ the value of an inner minimization problem over the matrix variable $X$. Writing $X=xx^{\top}+\xi$ and taking the SDP dual in the $\xi$ variable then results in the same saddle-point structure observed in Corollary 1. In other words, $\Gamma_{1}$ is the feasible domain to this partial dual of (5).
[…]
Figure 1: The sets ${\cal S}$, ${\cal S}_{\textup{SDP}}$, $\Gamma$, and $\Gamma^{\circ}$ are shown in blue, green, orange, and yellow respectively. By Remark 1, $(x,t)\in{\cal S}_{\textup{SDP}}$ if and only if $q(x)-2te_{\textup{obj}}\in\Gamma^{\circ}$.
Our results and analysis depend on the geometry of $\Gamma$, $\Gamma^{\circ}$, and their faces. We close this section with some properties of convex cones and their faces specialized to our setting. See [1] for more general statements.
Definition 2.
For a face ${\cal G}\trianglelefteq\Gamma^{\circ}$ and $(g_{\textup{obj}},g)\in\operatorname{rint}({\cal G})$, the conjugate face of ${\cal G}$ is
${\cal G}^{\triangle}\coloneqq\Gamma\cap{\cal G}^{\perp}=\Gamma\cap(g_{\textup{obj}},g)^{\perp}.$
Similarly, define the conjugate face of ${\cal F}$ for a face ${\cal F}\trianglelefteq\Gamma$.
Definition 3.
A face ${\cal G}\trianglelefteq\Gamma^{\circ}$ is exposed if there exists $(\gamma_{\textup{obj}},\gamma)\in\Gamma$ such that ${\cal G}=\Gamma^{\circ}\cap(\gamma_{\textup{obj}},\gamma)^{\perp}$.
We associate faces of $\Gamma$ and $\Gamma^{\circ}$ to points $(x,t)\in{\cal S}_{\textup{SDP}}$ as follows.
Definition 4.
Given $(x,t)\in{\cal S}_{\textup{SDP}}$, let ${\cal G}(x,t)\trianglelefteq\Gamma^{\circ}$ denote the minimal face of $\Gamma^{\circ}$ containing $q(x)-2te_{\textup{obj}}$ and define ${\cal F}(x,t)\coloneqq{\cal G}(x,t)^{\triangle}$.
The next fact follows from Definition 4.
Fact 1.
Given $(x,t)\in{\cal S}_{\textup{SDP}}$, we have that $q(x)-2te_{\textup{obj}}\in\operatorname{rint}({\cal G}(x,t))$ and ${\cal F}(x,t)=\Gamma\cap(q(x)-2te_{\textup{obj}})^{\perp}$.
3 Convex hull exactness
This section presents necessary and sufficient conditions for convex hull exactness, i.e., the property that $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$. The following result rephrases convex hull exactness as a question regarding the existence of nontrivial “rounding directions.” This result follows from basic convex analysis and is proved in Appendix A.
Lemma 1.
Suppose Assumption 1 holds. Then, $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ if and only if for every $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$, there exists a nonzero $(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}$ and $\epsilon>0$ such that $[(x,t)\pm\epsilon(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}.$
We capture the relevant set of “rounding directions” from Lemma 1 in the following definition.
Definition 5.
The subspace of rounding directions at $(x,t)\in{\cal S}_{\textup{SDP}}$ is
${\cal R}(x,t)\coloneqq\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\exists\epsilon>0\text{ s.t. }[(x,t)\pm\epsilon(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}\right\}.$
We say that this set is nontrivial if it contains a nonzero element.
Note that ${\cal R}(x,t)$ is a convex cone satisfying $-{\cal R}(x,t)={\cal R}(x,t)$ so that ${\cal R}(x,t)$ is indeed a subspace.
Remark 3.
One may compare our rounding directions to other similar definitions from elementary convex analysis [11, Section 5.1]. Fix a point $(x,t)\in{\cal S}_{\textup{SDP}}$ and $(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}$. Recall that $(x^{\prime},t^{\prime})$ is a feasible direction if there exists $\alpha>0$ such that $[(x,t),(x,t)+\alpha(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$. In particular, feasible directions are a unidirectional notion, whereas rounding directions are bidirectional. Next, recall that $(x^{\prime},t^{\prime})$ is a tangent direction if it is a limit of feasible directions. Again, tangent directions are unidirectional.
Remark 4.
Suppose $(x,t)\in{\cal S}_{\textup{SDP}}$ and $2t>\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Then, by Remark 1 there exists $\alpha>0$ such that $[(x,t)\pm\alpha(0_{n},1)]\subseteq{\cal S}_{\textup{SDP}}$. Thus, it suffices to verify the condition of Lemma 1 for points $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$ for which $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$.
3.1 Sufficient conditions for convex hull exactness
In general, the subspace of rounding directions ${\cal R}(x,t)$ is a difficult object to describe explicitly. Below, we identify a more manageable subset of ${\cal R}(x,t)$ that we will refer to as ${\cal R}^{\prime}(x,t)$. This will lead to a sufficient condition for convex hull exactness upon invoking Lemma 1.
Definition 6.
Let $(x,t)\in{\cal S}_{\textup{SDP}}$ and let ${\cal G}(x,t)$ be the minimal face of $\Gamma^{\circ}$ containing $q(x)-2te_{\textup{obj}}$. Define
${\cal R}^{\prime}(x,t)\coloneqq\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\operatorname{span}({\cal G}(x,t)),\,\forall\alpha\in{\mathbb{R}}\right\}.$
To motivate ${\cal R}^{\prime}(x,t)$, note that by Definitions 5 and 1, $(x^{\prime},t^{\prime})\in{\cal R}(x,t)$ if and only if $\exists\epsilon>0$ such that
$q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\Gamma^{\circ},\quad\forall\alpha\in[-\epsilon,\epsilon].$
If we assume (incorrectly) that $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}$ is linear in $\alpha$, then as ${\cal G}(x,t)$ is the minimal face of $\Gamma^{\circ}$ containing $q(x)-2te_{\textup{obj}}$, we have $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\operatorname{span}({\cal G}(x,t))$ for all $\alpha\in[-\epsilon,\epsilon]$ (in fact, also for all $\alpha\in{\mathbb{R}}$). Thus, $(x^{\prime},t^{\prime})\in{\cal R}^{\prime}(x,t)$. Unfortunately, $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}$ is nonlinear in $\alpha$ so that not all rounding directions from ${\cal R}(x,t)$ can be captured in this form. Below, we show that ${\cal R}^{\prime}(x,t)\subseteq{\cal R}(x,t)$. We will later give additional intuition and motivation for ${\cal R}^{\prime}(x,t)$ when $\Gamma^{\circ}$ is facially exposed.
Lemma 2.
Suppose Assumption 1 holds and $(x,t)\in{\cal S}_{\textup{SDP}}$. Then, ${\cal R}^{\prime}(x,t)\subseteq{\cal R}(x,t)$.
Proof.
Let $(x,t)\in{\cal S}_{\textup{SDP}}$ and $(x^{\prime},t^{\prime})\in{\cal R}^{\prime}(x,t)$. By continuity and the fact that $q(x)-2te_{\textup{obj}}\in\operatorname{rint}({\cal G}(x,t))$ (by Fact 1), there exists $\alpha>0$ such that for all $\epsilon\in[\pm\alpha]$ we have $q(x+\epsilon x^{\prime})-2(t+\epsilon t^{\prime})e_{\textup{obj}}\in{\cal G}(x,t)\subseteq\Gamma^{\circ}.$ By the third characterization of ${\cal S}_{\textup{SDP}}$ in Remark 1, we have that $[(x,t)\pm\alpha(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$. ∎
Lemmas 1 and 2 immediately imply the following sufficient condition for convex hull exactness.
Theorem 1.
Suppose Assumption 1 holds and that for all $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$, the set ${\cal R}^{\prime}(x,t)$ is nontrivial. Then, $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$.
The condition of Theorem 1 on the surface requires checking infinitely many systems of infinitely many constraints. Nonetheless, it is possible to check these conditions in some situations where $\Gamma$ (and related sets) possess simplifying symmetries. See Appendix B for some worked examples.
3.2 Necessary conditions for convex hull exactness
We now show that under a technical assumption (Assumption 2), we have ${\cal R}^{\prime}(x,t)={\cal R}(x,t)$. Thus, the sufficient condition of Theorem 1 is also necessary under the same assumption.
Assumption 2.
Suppose $\Gamma^{\circ}$ is facially exposed, i.e., every face of $\Gamma^{\circ}$ is exposed.
This assumption holds for any cone isomorphic to a slice of the nonnegative orthant, the second-order cone, or the positive semidefinite cone. See [19] for a longer discussion of this assumption and its connections to the nice cones. In general, all nice cones are facially exposed.
We will use the following property of exposed faces ${\cal G}\trianglelefteq\Gamma^{\circ}$.
Fact 2 ([1]).
[…]
We provide further intuition to the definition of ${\cal R}^{\prime}(x,t)$ under Assumption 2. Suppose $(x^{\prime},t^{\prime})\in{\cal R}(x,t)$. Noting that ${\cal F}(x,t)$ captures the convex quadratic inequalities defining ${\cal S}_{\textup{SDP}}$ that are tight at $(x,t)$, we deduce that $\left\langle(f_{\textup{obj}},f),q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\right\rangle$ must be zero for all $(f_{\textup{obj}},f)\in{\cal F}(x,t)$ and all small enough $\alpha$. In other words, a necessary condition for $(x^{\prime},t^{\prime})\in{\cal R}^{\prime}(x,t)$ is that
$q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\Gamma\cap{\cal F}(x,t)^{\perp}$
for all small enough $\alpha$. Under, Assumption 2 we have that $\Gamma\cap{\cal F}(x,t)^{\perp}={\cal G}(x,t)$. This same reasoning is presented formally in the lemma below.
Lemma 3.
Suppose Assumptions 1 and 2 hold and let $(x,t)\in{\cal S}_{\textup{SDP}}$. Then, ${\cal R}^{\prime}(x,t)={\cal R}(x,t)$.
Proof.
Fix $(x^{\prime},t^{\prime})\in{\cal R}(x,t)$. As ${\cal R}(x,t)$ is a convex cone, we may without loss of generality assume that $[(x,t)\pm(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$. Our goal is to show that $(x^{\prime},t^{\prime})\in{\cal R}^{\prime}(x,t)$, i.e., that
$q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\operatorname{span}({\cal G}(x,t)),\,\forall\alpha\in{\mathbb{R}}.$
As each coordinate of this vector is quadratic in $\alpha$, it suffices to show instead that the above inclusion holds for all $\alpha\in[-1,1]$. Indeed, suppose this inclusion has been shown for all $\alpha\in[-1,1]$ and let $\ell\in\operatorname{span}({\cal G}(x,t))^{\perp}$. Then, $\left\langle\ell,q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\right\rangle$ is a quadratic function in $\alpha$ mapping the interval $[-1,1]$ to 00. In particular, it must map all of ${\mathbb{R}}$ to zero. As $\ell\in\operatorname{span}({\cal G}(x,t))^{\perp}$ was arbitrary, we deduce that $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\operatorname{span}({\cal G}(x,t))$ for all $\alpha\in{\mathbb{R}}$.
Recall that ${\cal F}(x,t)=\Gamma\cap(q(x)-2te_{\textup{obj}})^{\perp}$ and let $(f_{\textup{obj}},f)\in\operatorname{rint}({\cal F}(x,t))$. By Assumptions 2 and 2, we may write ${\cal G}(x,t)=\Gamma^{\circ}\cap(f_{\textup{obj}},f)^{\perp}$. As $[(x,t)\pm(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$, we have that $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\Gamma^{\circ}$ for all $\alpha\in[-1,1]$. It remains to verify that the map
$\alpha\mapsto\left\langle(f_{\textup{obj}},f),q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\right\rangle$
evaluates to zero on $\alpha\in[-1,1]$. Again, as $[(x,t)\pm(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$, this map is nonpositive for all $\alpha\in[-1,1]$. Next, note that $(f_{\textup{obj}},f)\in{\cal F}(x,t)=\Gamma\cap(q(x)-2te_{\textup{obj}})^{\perp}$ so that this map evaluates to zero at $\alpha=0$. Finally, $(f_{\textup{obj}},f)\in\Gamma$ implies that this map is also convex. We conclude that this map is identically zero. ∎
The following necessary and sufficient condition for convex hull exactness then follows from Lemma 3.
Theorem 2.
Suppose Assumptions 1 and 2 hold. Then, $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ if and only if for all $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$, the set ${\cal R}^{\prime}(x,t)$ is nontrivial.
3.3 Explicit descriptions of ${\cal R}^{\prime}(x,t)$
Below, we give slightly simplified descriptions of ${\cal R}^{\prime}(x,t)$ (see Propositions 1 and 2) that may be useful in verifying the conditions of Theorems 1 and 2. See, for example, their use in Appendix B. The derivations of these descriptions require the following observation.
Observation 1.
Suppose Assumption 1 holds. Let $(x,t)\in{\cal S}_{\textup{SDP}}$ where $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Then, $\operatorname{span}({\cal G}(x,t))\not\supseteq{\mathbb{R}}\times 0_{m}$. In particular, ${\cal G}(x,t)^{\perp}=\operatorname{span}\left({\cal G}(x,t)^{\perp}\cap\left\{\gamma_{\textup{obj}}=1\right\}\right)$.
The proof of 1 and of Propositions 1 and 2 below are given in Appendix A.
Proposition 1.
Suppose Assumption 1 holds and let $(x,t)\in{\cal S}_{\textup{SDP}}$ where $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Then,
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}(x^{\prime})^{\top}A[\gamma]x^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\\ \left\langle A[\gamma]x+b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\end{array}\right\}.$
Proposition 2.
Suppose Assumptions 1 and 2 hold. Let $(x,t)\in{\cal S}_{\textup{SDP}}$ where $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$ and let $(1,f)\in\operatorname{rint}({\cal F}(x,t))$. Then,
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f]),\\ \left\langle A[\eta]x+b[\eta],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\eta)\in{\cal G}(x,t)^{\perp}\end{array}\right\}.$
3.4 Revisiting the setting of polyhedral $\Gamma$
Wang and Kılınç-Karzan [25] give sufficient conditions for convex hull exactness under the assumption that $\Gamma_{1}$ is polyhedral (equivalently, by Remark 1, whenever $\Gamma$ is polyhedral). This assumption holds, for example, when the set of quadratic forms $\left\{A_{\textup{obj}},A_{1},\dots,A_{m}\right\}$ is simultaneously diagonalizable. Specializing Theorem 2 to this setting, we prove the following necessary and sufficient counterpart to [25, Theorem 1].
{restatable}
theoremthmpolyhedral Suppose Assumption 1 holds and that $\Gamma$ is polyhedral. Then, for any $(x,t)\in{\cal S}_{\textup{SDP}}$ such that $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$ and any vector $f$ such that $(1,f)\in\operatorname{rint}({\cal F}(x,t))$, we have
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f]),\\ \left\langle b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal F}(x,t)\end{array}\right\}.$
Note that the only dependence on $(x,t)$ in the above description of ${\cal R}^{\prime}$ is in the definition of ${\cal F}={\cal F}(x,t)$. Hence, $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ if and only if the above set is nontrivial for every ${\cal F}\trianglelefteq\Gamma$ which is exposed by some vector $q(x)-2te_{\textup{obj}}$ for $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$.
Proof.
We begin by noting that when $\Gamma$ is polyhedral, so too is $\Gamma^{\circ}$ so that Assumption 2 holds. Next, we claim that for every face ${\cal G}\trianglelefteq\Gamma^{\circ}$ we have ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$. By definition, $\operatorname{span}({\cal G}^{\triangle})=\operatorname{span}(\Gamma\cap{\cal G}^{\perp})\subseteq{\cal G}^{\perp}$. On the other hand, as $\Gamma$ and $\Gamma^{\circ}$ are polyhedral, we have that $\dim({\cal G})+\dim({\cal G}^{\triangle})=m$ [22, Theorem 3]. Rearranging this equation, we have $\dim({\cal G}^{\triangle})=m-\dim({\cal G})=\dim({\cal G}^{\perp})$. We conclude that ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$.
Let $(x,t)\in{\cal S}_{\textup{SDP}}$ such that $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$ and let $(1,f)\in\operatorname{rint}({\cal F}(x,t))$. Then, 1 and Proposition 2 imply that
${\cal R}^{\prime}(x,t)$ =$=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f])\\ \left\langle A[\gamma]x+b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\end{array}\right\}$
=$=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f])\\ \left\langle A[\gamma]x+b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal F}(x,t)\end{array}\right\}$
=$=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f])\\ \left\langle b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal F}(x,t)\end{array}\right\}.$
Here, the second line follows because we have shown ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$ holds for every face ${\cal G}\trianglelefteq\Gamma^{\circ}$ and by definition ${\cal F}(x,t)={\cal G}(x,t)^{\triangle}$. The third line follows from the fact that $(1,f)\in\operatorname{rint}({\cal F}(x,t))$ implies $\ker(A[f])\subseteq\ker(A[\gamma])$ for every $(1,\gamma)\in{\cal F}(x,t)$. The result then follows from Theorem 2. ∎
Remark 5.
The main difference between Section 3.4 and [25, Theorem 1] is that Section 3.4 only considers certain (a fortiori semidefinite) faces of $\Gamma$ whereas [25, Theorem 1] imposes a constraint on every semidefinite face of $\Gamma$. This idea of restricting the analysis to certain faces of $\Gamma$ was previously investigated by [6, 17] who used it to provide sufficient conditions for objective value exactness.
Section 3.4 can be used to characterize when the SDP relaxation of a set defined by two quadratic constraints is exact; see Appendix B. This complements work of [26] that describes the convex hull of a set defined by two quadratic constraints in all settings but requires a strengthened SDP relaxation.
Acknowledgments
This work was supported by the National Science Foundation [grant number CMMI 1454548], the Office of Naval Research [grant number N00014-19-1-2321], and the Air Force Office of Scientific Research [grant number FA95502210365].
Appendix A Deferred proofs
Proof of Remark 1.
Fix $(x,t)\in{\mathbb{R}}^{n+1}$. Note that
$\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$ $=\sup_{\gamma\in{\mathbb{R}}_{+}^{m}}\left\{[\gamma,q(x)]:\,A[\gamma]\succeq 0\right\}$
$=\inf_{\xi\in{\mathbb{S}}^{n}_{+}}\left\{q_{\textup{obj}}(x)+\left\langle A_{\textup{obj}},\xi\right\rangle:\,q_{i}(x)+\left\langle A_{i},\xi\right\rangle\leq 0,\forall i\in[m]\right\}.$
Here, the second equality follows from the strong conic duality theorem and Assumption 1. Taking $X\coloneqq xx^{\top}+\xi$, we deduce that the first equality in Remark 1 holds.
By Assumption 1, $\Gamma=\operatorname{clcone}\left\{(1,\gamma):\,\gamma\in\Gamma_{1}\right\}$ so that $[\gamma,q(x)]\leq 2t$ for all $\gamma\in\Gamma_{1}$ if and only if $\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle\leq 2\gamma_{\textup{obj}}t$ for all $(\gamma_{\textup{obj}},\gamma)\in\Gamma$; this gives the second equality. The third equality holds by definition of the polar cone. ∎
Proof of Lemma 1.
Note that ${\cal S}_{\textup{SDP}}$ is a closed convex set containing no lines. It is clear that $(0_{n},1)$ is a recessive direction of ${\cal S}_{\textup{SDP}}$. We claim that $(0_{n},1)$ is the only recessive direction: Let $\gamma^{*}$ such that $A[\gamma^{*}]\succ 0$ (this is possible by Assumption 1) and consider any $(x^{\prime},t^{\prime})$ where $x^{\prime}$ is nonzero. Then, for any $(\tilde{x},\tilde{t})\in{\cal S}_{\textup{SDP}}$, we have $2(\tilde{t}+\alpha t^{\prime})<[\gamma^{*},q(\tilde{x}+\alpha x^{\prime})]$ for all $\alpha>0$ large enough. We deduce by [20, Theorem 18.5], that ${\cal S}_{\textup{SDP}}$ is the sum of the convex hull of its extreme points and the direction $(0_{n},1)$. In particular $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ if and only if $(x,t)$ is not extreme for every $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. By definition, $(x,t)$ is not extreme if and only if there exists a nonzero $(x^{\prime},t^{\prime})$ and $\alpha>0$ such that $[(x,t)\pm\alpha(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$. ∎
Proof of 1.
Suppose $\operatorname{span}({\cal G}(x,t))\supseteq{\mathbb{R}}\times 0_{m}$ so that $(0_{n},1)\in{\cal R}^{\prime}(x,t)$. By Lemma 2, there exists $\alpha>0$ such that $(x,t-\alpha)\in{\cal S}_{\textup{SDP}}$. This contradicts $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Thus, we deduce that $\operatorname{span}({\cal G}(x,t))\not\supseteq{\mathbb{R}}\times 0_{m}$. Equivalently, ${\cal G}(x,t)^{\perp}\not\subseteq 0\times{\mathbb{R}}^{m}$ and there exists $(1,\bar{\gamma})\in{\cal G}(x,t)^{\perp}$. Then, for any $(\gamma_{\textup{obj}},\gamma)\in{\cal G}(x,t)^{\perp}$, we can write $(\gamma_{\textup{obj}},\gamma)$ as a linear combination of $(\gamma_{\textup{obj}},\gamma)+(1-\gamma_{\textup{obj}})(1,\bar{\gamma})$ and $(1,\bar{\gamma})$. ∎
Proof of Proposition 1.
Note that $(x^{\prime},t^{\prime})\in{\cal R}^{\prime}(x,t)$ if and only if for all $(\gamma_{\textup{obj}},\gamma)\in{\cal G}(x,t)^{\perp}$, the expression
$\left\langle(\gamma_{\textup{obj}},\gamma),q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\right\rangle$
$\quad=\alpha^{2}(x^{\prime})^{\top}A(\gamma_{\textup{obj}},\gamma)(x^{\prime})$
$\qquad+2\alpha\left(\left\langle A(\gamma_{\textup{obj}},\gamma)x+b(\gamma_{\textup{obj}},\gamma),x^{\prime}\right\rangle-t^{\prime}\right)$
$\qquad+\left\langle(\gamma_{\textup{obj}},\gamma),q(x)-2te_{\textup{obj}}\right\rangle$
is identically zero in $\alpha$. This occurs if and only if for all $(\gamma_{\textup{obj}},\gamma)\in{\cal G}(x,t)^{\perp}$,
$(x^{\prime})^{\top}A(\gamma_{\textup{obj}},\gamma)x^{\prime}=0,\quad\text{and}$
$\left\langle A(\gamma_{\textup{obj}},\gamma)x+b(\gamma_{\textup{obj}},\gamma),x^{\prime}\right\rangle-t^{\prime}=0.$
The result then follows from 1. ∎
Proof of Proposition 2.
Let $(1,f)\in\operatorname{rint}({\cal F}(x,t))$. By Proposition 1, it suffices to show that $x^{\prime}\in\ker(A[f])$ if and only if
$(x^{\prime})^{\top}A[\gamma]x^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}.$
The reverse direction holds immediately as $(1,f)\in{\cal F}(x,t)\subseteq{\cal G}(x,t)^{\perp}$ and $A[f]\succeq 0$. To see the forward direction: Let $x^{\prime}\in\ker(A[f])$ and set $v_{\textup{obj}}=(x^{\prime})^{\top}A_{\textup{obj}}x^{\prime}$. Similarly, set $v_{i}=(x^{\prime})^{\top}A_{i}x^{\prime}$. Then, $\forall(\gamma_{\textup{obj}},\gamma)\in\Gamma$, we have
$\left\langle(v_{\textup{obj}},v),(\gamma_{\textup{obj}},\gamma)\right\rangle$ $=(x^{\prime})^{\top}A(\gamma_{\textup{obj}},\gamma)x^{\prime}\geq 0.$
Thus, $(-v_{\textup{obj}},-v)\in\Gamma^{\circ}$. On the other hand, $\left\langle(v_{\textup{obj}},v),(1,f)\right\rangle=(x^{\prime})^{\top}A[f]x^{\prime}=0$. We deduce that $(-v_{\textup{obj}},-v)\in\Gamma^{\circ}\cap(1,f)^{\perp}={\cal F}(x,t)^{\triangle}={\cal G}(x,t)$. In particular, $(x^{\prime})^{\top}A(\gamma_{\textup{obj}},\gamma)x^{\prime}=\left\langle(v_{\textup{obj}},v),(\gamma_{\textup{obj}},\gamma)\right\rangle=0$ for all $(\gamma_{\textup{obj}},\gamma)\in{\cal G}(x,t)^{\perp}$. ∎
Appendix B Example applications of Theorems 1, 2 and 3.4
In this section, we apply the results of Section 3 to three problems. These examples show that the conditions of Theorems 1, 2 and 3.4 can be checked in situations where $\Gamma$ is sufficiently simple.
B.1 Mixed binary programming
To begin, we apply our results to a well-studied prototypical set involving a convex quadratic function, a binary variable and a big-M relation. This set occurs as a substructure commonly studied in sparse regression applications. The convex hull characterization of this set is well-known in the literature and is often shown as a consequence of the perspective formulation trick due to Ceria and Soares [7]. The example in this subsection highlights the systematic nature of our approach.
Consider the epigraph set
${\cal S}=\left\{(x,t)\in{\mathbb{R}}^{2}\times{\mathbb{R}}:\,\begin{array}[]{l}q_{\textup{obj}}(x)\coloneqq x_{2}^{2}\leq 2t\\ q_{1}(x)\coloneqq x_{1}(x_{1}-1)=0\\ q_{2}(x)\coloneqq\sqrt{2}x_{2}(x_{1}-1)=0\end{array}\right\}.$
In words, $x_{1}$ is a binary on-off variable, $x_{2}$ is a continuous variable which is constrained to be off whenever $x_{1}$ is off, and $t$ is the epigraph variable corresponding to $x_{2}^{2}$. The normalization of $q_{2}(x)$ is not important here and is made only for notational convenience in the calculations. Additionally, the presence of equality constraints does not markedly affect our results; the only change in this case is that the corresponding $\gamma$ multiplier is no longer constrained to be nonnegative.
It is well-known that $\operatorname{conv}({\cal S})$ is given by the perspective reformulation of ${\cal S}$ (see e.g., [9]), i.e.,
[…]
We give an alternative proof of (6). We will show that $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$, the projected SDP relaxation, using Theorem 2. Then, using an explicit description of $\Gamma^{\circ}$, we will give a description of $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ in the original space.
A simple computation shows that in this setting, we have
$\Gamma=\left\{(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}^{3}:\,\gamma_{\textup{obj}}+\gamma_{1}\geq\sqrt{(\gamma_{\textup{obj}}-\gamma_{1})^{2}+\left(\sqrt{2}\gamma_{2}\right)^{2}}\right\}\text{ and}$
$\Gamma^{\circ}=\left\{(\ell_{\textup{obj}},\ell)\in{\mathbb{R}}^{3}:\,-\ell_{\textup{obj}}-\ell_{1}\geq\sqrt{(\ell_{\textup{obj}}-\ell_{1})^{2}+\left(\sqrt{2}\ell_{2}\right)^{2}}\right\}.$
In words, $\Gamma$ and $\Gamma^{\circ}$ are both (rotated) second-order cones and Assumptions 1 and 2 hold.
It remains to show that for all $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$, the set ${\cal R}^{\prime}(x,t)$ is nontrivial. To this end, let $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. Note that as the point $(0,0)\in{\cal S}$, we deduce $(x,t)\neq(0,0)$. Recall that $\Gamma^{\circ}$ has three types of faces: the two trivial faces (the apex and the cone itself) and the one-dimensional proper faces. Thus, there are three cases to consider: (i) ${\cal G}(x,t)=\left\{0\right\}$, (ii) ${\cal G}(x,t)=\Gamma^{\circ}$, and (iii) ${\cal G}(x,t)$ is a one-dimensional face of $\Gamma^{\circ}$.
In case (i), $q(x)-2te_{\textup{obj}}=0$ implying that $(x,t)\in{\cal S}$, a contradiction. In case (ii), $\operatorname{span}({\cal G}(x,t))={\mathbb{R}}^{3}$ so that ${\cal R}^{\prime}(x,t)={\mathbb{R}}^{3}$ and is nontrivial. In the final case, a mechanical application of Proposition 2 gives
${\cal R}^{\prime}(x,t)$ =$=\left\{\begin{array}[]{r}\begin{pmatrix}2t\\ -x_{2}\\ 0\end{pmatrix},\,\begin{pmatrix}-x_{2}\\ x_{1}\\ 0\end{pmatrix},\,\begin{pmatrix}t\\ -x_{2}\\ x_{1}\end{pmatrix},\,\begin{pmatrix}x_{2}(x_{1}-1+2t)\\ -x_{1}^{2}+x_{1}-2tx_{1}-2t\\ 2x_{2}\end{pmatrix}\end{array}\right\}^{\perp}.$ (20)
In this case, one may verify that $(x,t)\in{\cal R}^{\prime}(x,t)$. Then, as $(x,t)$ is nonzero we deduce that ${\cal R}^{\prime}(x,t)$ is nontrivial in this case as well.
As ${\cal R}^{\prime}(x,t)$ is nontrivial in all three cases, we conclude from Theorem 2 that
$\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}=\left\{\begin{array}[]{l}(x,t)\in{\mathbb{R}}^{3}:\\ \leavevmode\nobreak\ \leavevmode\nobreak\ -(q_{\textup{obj}}(x)-2t)-q_{1}(x)\\ \qquad\geq\sqrt{(q_{\textup{obj}}(x)-2t-q_{1}(x))^{2}+2q_{2}(x)^{2}}\end{array}\right\}.$
Here, the second equation follows from the last equivalence of ${\cal S}_{\textup{SDP}}$ in Remark 1 and the characterization of $\Gamma^{\circ}$ derived as the rotated second-order cone for this specific set ${\cal S}$. Substituting back the definitions of $q_{\textup{obj}},q_{1},q_{2}$ and squaring both sides of the constraint in the set ${\cal S}_{\textup{SDP}}$ above gives
$\operatorname{conv}({\cal S})=\left\{(x,t)\in{\mathbb{R}}^{3}:\,\begin{array}[]{l}-(x_{2}^{2}-2t)-x_{1}(x_{1}-1)\geq 0\\ 4(1-x_{1})(2tx_{1}-x_{2}^{2})\geq 0\end{array}\right\}.$
While the two descriptions of $\operatorname{conv}({\cal S})$ given in (6) and given above are different, we can check that they indeed define the same set. First, the $\subseteq$ inclusion follows immediately as $x_{2}^{2}-2t\leq x^{2}-2tx_{1}\leq 0$ so that both constraints in the description above are satisfied. In the reverse direction, we first establish that every $(x,t)$ in the set above satisfies $x_{1}\in[0,1]$. If $x_{1}<0$, then the first constraint implies $t>0$ while the second constraint implies $t\leq 0$, a contradiction. If $x_{1}>1$, then the second constraint implies $2t\leq x_{2}^{2}/x_{1}$. On the other hand, the first constraint implies $0<x_{1}(1-x_{1})\leq-(x_{2}^{2}-2t)\leq x_{2}^{2}(1/x_{1}-1)$, which contradicts $x_{1}>1$. Finally, suppose $(x,t)$ is in the set above with $x_{1}\in[0,1]$. We will show that $(x,t)$ satisfies the constraints in (6). If $x_{1}=1$, then $x_{2}^{2}-2tx_{1}=x_{2}^{2}-2t\leq 0$ by the first constraint (whence also $t\geq 0$). If $x_{1}\in(0,1)$, then the second constraint is equivalent to $x_{2}^{2}-2tx_{1}\leq 0$ (again, $t\geq 0$). Finally, if $x_{1}=0$, then we have $x_{2}^{2}-2tx_{1}\leq 0$ by the second constraint and $x_{2}^{2}-2t\leq 0$ by the first constraint (so that $t\geq 0$).
B.2 Quadratic matrix programs
Quadratic matrix programs (QMPs) [2, 25] are a generalization of QCQPs where the decision variable $x\in{\mathbb{R}}^{n}$ is replaced by a decision matrix $X\in{\mathbb{R}}^{r\times k}$. These problems find a variety of application and have been used to model robust least squares problems, the orthogonal Procrustes problem, and certain sphere packing problems. Formally, a QMP is an optimization problem in the variable $X\in{\mathbb{R}}^{r\times k}$, where the constraints and objective function are each of the form
$\operatorname{tr}(X^{\top}\mathbb{A}X)+2\operatorname{tr}(B^{\top}X)+c$
for some $\mathbb{A}\in{\mathbb{S}}^{r}$, $B\in{\mathbb{R}}^{r\times k}$, and $c\in{\mathbb{R}}$.
Alternatively, letting $x\in{\mathbb{R}}^{n}$ (resp. $b\in{\mathbb{R}}^{n}$) denote the vector formed by stacking the columns of $X$ (resp. $B$) on top of each other, we can rewrite the above expression as
$x^{\top}(I_{k}\otimes\mathbb{A})x+2\left\langle b,x\right\rangle+c.$
We will choose to view QMPs as the special class of QCQPs where the quadratic forms $A_{\textup{obj}},A_{1},\dots,A_{m}$ are each of the form $I_{k}\otimes\mathbb{A}$ for some $\mathbb{A}\in{\mathbb{S}}^{r}$.
We will next establish that if the number of constraints is small compared to $k$ (originally the width of the matrix variable), then convex hull exactness holds. Specifically, we show below that if Assumption 1 holds, $A_{\textup{obj}}=I_{k}\otimes\mathbb{A}_{\textup{obj}}$, $A_{1}=I_{k}\otimes\mathbb{A}_{1}$, …, $A_{m}=I_{k}\otimes\mathbb{A}_{m}$ for some $\mathbb{A}_{\textup{obj}},\mathbb{A}_{1},\dots,\mathbb{A}_{m}\in{\mathbb{S}}^{r}$ and $k\geq m$, then ${\cal R}^{\prime}(x,t)$ is nontrivial for every $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. By Theorem 1, this will imply that convex hull exactness holds, i.e., $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$.
Fix $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. Based on Theorems 1 and 1, our goal is to prove that
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime\top}A[\gamma]x^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\\ \left\langle A[\gamma]x+b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\end{array}\right\}$ (23)
is nontrivial. We claim that it suffices to show how to construct a nonzero $y\in{\mathbb{R}}^{r}$ such that
$y^{\top}\mathbb{A}[\gamma]y=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}.$ (24)
To see that this suffices, note that for any $w\in{\mathbb{R}}^{k}$, the vector $x^{\prime}\coloneqq w\otimes y$ satisfies the first constraint in (23) since for $(1,\gamma)\in{\cal G}(x,t)^{\perp}$, we have
$(w\otimes y)^{\top}A[\gamma](w\otimes y)=(w^{\top}w)(y^{\top}\mathbb{A}[\gamma]y)=0.$
Then, $(w\otimes y,t^{\prime})\in{\cal R}^{\prime}(x,t)$ if and only if the following relation also holds
$\left\langle A[\gamma]x+b[\gamma],w\otimes y\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}.$
This is a system of $\dim({\cal G}(x,t)^{\perp})$-many homogeneous linear equations in the variables $(w,t^{\prime})\in{\mathbb{R}}^{k+1}$. Note that as ${\cal G}(x,t)\ni q(x)-2te_{\textup{obj}}$, which is nonzero by assumption, we have that $\dim({\cal G}(x,t)^{\perp})\leq m$. As $k+1>m$ by assumption, we deduce that this system has a nontrivial solution. Thus, we conclude that (23) is nontrivial if there exists a nonzero $y\in{\mathbb{R}}^{r}$ satisfying (24).
It remains to construct $y$. By definition of ${\cal S}_{\textup{SDP}}$, there exists $Y\succeq 0$ such that
$\begin{cases}q_{\textup{obj}}(x)+\left\langle A_{\textup{obj}},Y\right\rangle\leq 2t,\\ q_{i}(x)+\left\langle A_{i},Y\right\rangle\leq 0,\,\forall i\in[m].\end{cases}$ (25)
Without loss of generality, we can assume $Y=({1\over k}I_{k})\otimes\mathbb{Y}$. As $(x,t)\notin{\cal S}$, we have that $\mathbb{Y}\in{\mathbb{S}}^{r}_{+}\setminus\left\{0\right\}$ and so we may pick a nonzero $y\in{\mathbb{R}}^{r}$ such that $yy^{\top}\preceq\mathbb{Y}$. For notational convenience, let $\ell_{\textup{obj}}\coloneqq y^{\top}\mathbb{A}_{\textup{obj}}y$ and $\ell_{i}\coloneqq y^{\top}\mathbb{A}_{i}y$ for $i\in[m]$. Note that for any $(\gamma_{\textup{obj}},\gamma)\in\Gamma$, we have $A(\gamma_{\textup{obj}},\gamma)\succeq 0$, or equivalently $\mathbb{A}(\gamma_{\textup{obj}},\gamma)\succeq 0$. Thus, $yy^{\top}\preceq\mathbb{Y}$ implies that $\left\langle(\gamma_{\textup{obj}},\gamma),(\ell_{\textup{obj}},\ell)\right\rangle=y^{\top}\mathbb{A}(\gamma_{\textup{obj}},\gamma)y\leq\left\langle\mathbb{A}(\gamma_{\textup{obj}},\gamma),\mathbb{Y}\right\rangle$. Also, from $Y=({1\over k}I_{k})\otimes\mathbb{Y}$ and the relation between the matrices $A_{\textup{obj}},A_{i}$ and $\mathbb{A}_{\textup{obj}},\mathbb{A}_{i}$, we have $\left\langle\mathbb{A}(\gamma_{\textup{obj}},\gamma),\mathbb{Y}\right\rangle=\left\langle A(\gamma_{\textup{obj}},\gamma),Y\right\rangle$. We deduce that for $(\gamma_{\textup{obj}},\gamma)\in\Gamma$,
$\left\langle\begin{pmatrix}\gamma_{\textup{obj}}\\ \gamma\end{pmatrix},q(x)-2te_{\textup{obj}}+\begin{pmatrix}\ell_{\textup{obj}}\\ \ell\end{pmatrix}\right\rangle\leq\left\langle\begin{pmatrix}\gamma_{\textup{obj}}\\ \gamma\end{pmatrix},q(x)-2te_{\textup{obj}}+\begin{pmatrix}\left\langle A_{\textup{obj}},Y\right\rangle\\ \left(\left\langle A_{i},Y\right\rangle\right)_{i}\end{pmatrix}\right\rangle\leq 0,$
where the last inequality follows from (25) and $(\gamma_{\textup{obj}},\gamma)\in\Gamma\subseteq{\mathbb{R}}^{1+m}_{+}$. This then shows that $q(x)-2te_{\textup{obj}}+(\ell_{\textup{obj}},\ell)\in\Gamma^{\circ}$. Moreover, because $\mathbb{A}(\gamma_{\textup{obj}},\gamma)\succeq 0$, we have
$0\geq-y^{\top}\mathbb{A}(\gamma_{\textup{obj}},\gamma)y=\left\langle\begin{pmatrix}\gamma_{\textup{obj}}\\ \gamma\end{pmatrix},\begin{pmatrix}-\ell_{\textup{obj}}\\ -\ell\end{pmatrix}\right\rangle,$
which implies $-(\ell_{\textup{obj}},\ell)\in\Gamma^{\circ}$. We have shown that $q(x)-2te_{\textup{obj}}+(\ell_{\textup{obj}},\ell)$ and $-(\ell_{\textup{obj}},\ell)$ both lie in $\Gamma^{\circ}$. Then, as $q(x)-2te_{\textup{obj}}\in\operatorname{rint}({\cal G}(x,t))$, we deduce that $(\ell_{\textup{obj}},\ell)\in\operatorname{span}({\cal G}(x,t))$. In particular, $y^{\top}\mathbb{A}(\gamma_{\textup{obj}},\gamma)y=\left\langle(\gamma_{\textup{obj}},\gamma),(\ell_{\textup{obj}},\ell)\right\rangle=0$ for all $(\gamma_{\textup{obj}},\gamma)\in{\cal G}(x,t)^{\perp}$.
Remark 6.
SDP exactness in the context of QMPs was previously studied by Wang and Kılınç-Karzan [25], Beck [2]. Specifically, under Assumption 1, Beck [2] shows that objective value exactness holds whenever $k\geq m$ and Wang and Kılınç-Karzan [25] show that convex hull exactness holds whenever $k\geq m+2$. The above calculation strengthens both of these results by showing that under Assumption 1 convex hull exactness holds whenever $k\geq m$.
B.3 Sets defined by two quadratic constraints
We apply our results to derive a characterization of convex hull exactness of SDP relaxation when we have only two quadratic constraints. Note that when such a set is obtained from the epigraphical representation of a quadratic objective function and a single quadratic constraint, SDP convex hull exactness is well-studied, see e.g., [24] and references therein. In contrast, we will show below that the case of two general quadratic constraints is more nuanced. In the case of strict inequalities and under a strict feasibility assumption, Yıldıran [26] establishes that the convex hull of a set defined by two quadratic constraints is given by two constraints admitting second-order cone (SOC) representations and these constraints can be obtained by examining the eigenvalues of certain Lagrange aggregation matrices. Our results will present a simple way of checking exactly when such a convex hull characterization is given by its SDP relaxation as opposed to the more involved condition of [26].
Consider a set defined by two quadratic constraints, $\hat{\cal S}\coloneqq\left\{x\in{\mathbb{R}}^{n}:\,q_{1}(x)\leq 0,q_{2}(x)\leq 0\right\}$. To connect such a set to the epigraphical sets studied in the main body, we will define $q_{\textup{obj}}(x)\coloneqq 0$. Then, Assumption 1 is equivalent to the assumption that there exists $\hat{\gamma}=(\hat{\gamma}_{1},\hat{\gamma}_{2})\in{\mathbb{R}}^{2}_{+}$ such that $\hat{\gamma}_{1}A_{1}+\hat{\gamma}_{2}A_{2}\succ 0$. The projected SDP relaxation of this set is given by
$\hat{\cal S}_{\textup{SDP}}\coloneqq\left\{x\in{\mathbb{R}}^{n}:\,\begin{array}[]{l}\exists X\succeq xx^{\intercal}\\ \left\langle A_{i},X\right\rangle+2b_{i}^{\intercal}x+c_{i}\leq 0,\forall i=1,2\end{array}\right\}.$
Our goal is to understand when it holds that $\operatorname{conv}(\hat{\cal S})=\hat{\cal S}_{\textup{SDP}}$. Recognizing these sets as the $x$-components of ${\cal S}$ and ${\cal S}_{\textup{SDP}}$, we deduce the equivalence holds if and only if $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$.
As $A_{\textup{obj}}=0$, we have that
$\Gamma=\left\{(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}^{1+2}_{+}:\,\gamma_{1}A_{1}+\gamma_{2}A_{2}\succeq 0\right\}.$
Note that $\Gamma_{1}$ is a cone in ${\mathbb{R}}^{2}$ so that it is necessarily polyhedral. We deduce that $\Gamma$ is also polyhedral and we may apply Section 3.4 in this situation. Note that as $A[\hat{\gamma}]\succ 0$, we must have that $\Gamma_{1}$ is full dimensional. Let $\gamma^{(1)},\gamma^{(2)}\in{\mathbb{R}}^{2}_{+}$ be generators of $\Gamma_{1}$, i.e., such that $\Gamma_{1}=\operatorname{cone}(\left\{\gamma^{(1)},\gamma^{(2)}\right\})$. Then, $\Gamma=\operatorname{cone}\left(\left\{\begin{pmatrix}0\\ \gamma^{(1)}\end{pmatrix},\begin{pmatrix}0\\ \gamma^{(2)}\end{pmatrix},e_{\textup{obj}}\right\}\right)$ so that $\Gamma$ has exactly 6 proper faces that we deal with individually below.
Suppose ${\cal F}=\operatorname{cone}(e_{\textup{obj}})$. Then,
$\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(0)\\ -t^{\prime}=0\end{array}\right\}\supseteq{\mathbb{R}}^{n}\times\left\{0\right\}$
is nontrivial.
Next, suppose ${\cal F}=\operatorname{cone}((0,\gamma^{(i)}))$ for some $i=1,2$ or ${\cal F}=\operatorname{cone}(\left\{(0,\gamma^{(1)}),(0,\gamma^{(2)})\right\})$. Note that ${\cal F}$ does not contain any points of the form $(1,f)$. Thus, by 1, ${\cal F}$ is not exposed by any vector $q(x)-2te_{\textup{obj}}$ where $x,t$ satisfy $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Section 3.4 then places no constraint on the corresponding ${\cal R}^{\prime}$.
Finally, suppose ${\cal F}=\operatorname{cone}(\left\{e_{\textup{obj}},(0,\gamma^{(1)})\right\})$. The case $\gamma^{(2)}$ is analogous. Note that $(1,\gamma^{(1)})\in\operatorname{rint}({\cal F})$. The relevant set in Section 3.4 is
$\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[\gamma^{(1)}])\\ \left\langle b[\gamma^{(1)}],x^{\prime}\right\rangle=0\\ t^{\prime}=0\end{array}\right\}.$
This set is nontrivial if and only if $\ker(A[\gamma^{(1)}])\cap b[\gamma^{(1)}]^{\perp}$ is nontrivial. Section 3.4 requires that this set is nontrivial if ${\cal F}$ is exposed by some vector of the form $q(x)-2te_{\textup{obj}}$ for some $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$, i.e., if $x\in\hat{{\cal S}}_{\textup{SDP}}\setminus\hat{{\cal S}}$ and $t$ satisfy $t=0$ and
$\begin{cases}[\gamma^{(1)},q(x)]=0,\\ [\gamma^{(2)},q(x)]<0.\end{cases}$
Suppose this system is feasible and $x$ satisfies the above relations. Our goal is to show that $\gamma^{(1)}\in{\mathbb{R}}^{2}_{++}$. Suppose for the sake of contradiction that $\gamma^{(1)}=e_{1}$. The above relations would imply that $0=[\gamma^{(1)},q(x)]=q_{1}(x)$ and $0>[\gamma^{(2)},q(x)]=\gamma^{(2)}_{1}q_{1}(x)+\gamma^{(2)}_{2}q_{2}(x)=\gamma^{(2)}_{2}q_{2}(x)$. We see that $\gamma^{(2)}_{2}\neq 0$, and in fact $\gamma^{(2)}_{2}>0$ (as $\gamma^{(2)}\in{\mathbb{R}}^{2}_{+}$), and thus $0>q_{2}(x)$. Hence, $x\in\hat{{\cal S}}$, a contradiction. In particular, if $\gamma^{(1)}=e_{1}$, then ${\cal F}$ is not exposed by some vector of the form $q(x)-2te_{\textup{obj}}$ with $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. Similarly, if $\gamma^{(1)}=e_{2}$, then ${\cal F}$ is not exposed by some vector of the form $q(x)-2te_{\textup{obj}}$ with $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. Note also that if $\gamma^{(1)}\in{\mathbb{R}}^{2}_{++}$ and $[\gamma^{(1)},q(x)]=0>[\gamma^{(2)},q(x)]$, then $x\in\hat{\cal S}_{\textup{SDP}}\setminus\hat{\cal S}$.
We collect the above discussion in the following lemma.
Lemma 4.
Suppose there exists $\hat{\gamma}\in{\mathbb{R}}^{2}_{+}$ such that $A[\hat{\gamma}]\succ 0$. Let $\gamma^{(1)},\gamma^{(2)}\in{\mathbb{R}}^{2}_{+}$ be generators of $\Gamma_{1}$. If $\gamma^{(1)}\in{\mathbb{R}}^{2}_{++}$ and there exists $x\in{\mathbb{R}}^{n}$ such that $[\gamma^{(1)},q(x)]=0$ and $[\gamma^{(2)},q(x)]<0$, then we require $\ker(A[\gamma^{(1)}])\cap b[\gamma^{(1)}]^{\perp}$ is nontrivial. Similarly, if $\gamma^{(2)}\in{\mathbb{R}}^{2}_{++}$ and there exists $x\in{\mathbb{R}}^{n}$ such that $[\gamma^{(2)},q(x)]=0$ and $[\gamma^{(1)},q(x)]<0$, then we require $\ker(A[\gamma^{(2)}])\cap b[\gamma^{(2)}]^{\perp}$ is nontrivial. Then, $\operatorname{conv}(\hat{\cal S})=\hat{\cal S}_{\textup{SDP}}$ if and only if the requirements stated above hold.
In particular, whenever we have $q_{1}$ and $q_{2}$ are both nonconvex, primal strict feasibility is satisfied (i.e., $\exists\hat{x}\in{\mathbb{R}}^{n}$ such that $q_{1}(\hat{x}),q_{2}(\hat{x})<0$) and dual strict feasibility is satisfied (i.e., $\exists\hat{\gamma}\in{\mathbb{R}}^{2}_{+}$ such that $A[\hat{\gamma}]\succ 0$), then $\operatorname{conv}(\hat{\cal S})=\hat{\cal S}_{\textup{SDP}}$ if and only if for both $i=1,2$, we have that
$\ker(A[\gamma^{(i)}])\cap b[\gamma^{(i)}]^{\perp}\text{ is nontrivial}.$
Proof.
The first claim has already been proved in the preceding discussion under the dual strict feasibility assumption. The remainder of the proof further assumes that $q_{1}$ and $q_{2}$ are both nonconvex and that there exists $\hat{x}\in{\mathbb{R}}^{n}$ such that $q_{1}(\hat{x}),q_{2}(\hat{x})<0$. Under these additional assumptions, we will show that $\gamma^{(1)}\in{\mathbb{R}}^{2}_{++}$ and there exists $x\in{\mathbb{R}}^{n}$ such that $[\gamma^{(1)},q(x)]=0$ and $[\gamma^{(2)},q(x)]<0$. An identical proof will show that the analogous statement with the roles of $\gamma^{(1)}$ and $\gamma^{(2)}$ interchanged holds. This will complete the proof of the second claim.
First, $\gamma^{(1)},\gamma^{(2)}\in{\mathbb{R}}^{2}_{++}$ follows from the assumption that $q_{1},\,q_{2}$ are both nonconvex. As $\gamma^{(i)}\in{\mathbb{R}}^{2}_{++}$, we deduce that $A[\gamma^{(i)}]$ has a nontrivial kernel for both $i=1,2$. Note that $[\gamma^{(i)},q(\hat{x})]<0$ for both $i=1,2$. Now pick a nonzero vector $y\in\ker(A[\gamma^{(2)}])$. Then,
$[\gamma^{(2)},q(\hat{x}+\alpha y)]=q(\hat{x})+2\alpha b[\gamma^{(2)}]^{\top}y$
is a linear function in $\alpha$ that is negative at $\alpha=0$ and that is without loss of generality nonincreasing as $\alpha\to\infty$ (otherwise, negate $y$). On the other hand, $[\hat{\gamma},q(\hat{x}+\alpha y)]$ is a strongly convex quadratic function in $\alpha$ so that it is positive for some $\alpha$ large enough. As $\hat{\gamma}\in\Gamma_{1}=\operatorname{cone}(\left\{\gamma^{(1)},\gamma^{(2)}\right\})$, we deduce that $[\gamma^{(1)},q(\hat{x}+\alpha y)]$ is positive for all $\alpha>0$ large enough. As $[\gamma^{(1)},q(\hat{x}+\alpha y)]$ is a continuous function of $\alpha$ and $[\gamma^{(1)},q(\hat{x})]<0$, there exists some $\alpha>0$ for which $[\gamma^{(1)},q(\hat{x}+\alpha y)]=0$ and $[\gamma^{(2)},q(\hat{x}+\alpha y)]<0$. ∎
Remark 7.
Lemma 4 presents an easy to check condition for SDP convex hull exactness in the case of two quadratic constraints. One may compare this to the main result in [26] which gives a description of $\operatorname{clconv}(\hat{\cal S})$ assuming that $\hat{\cal S}$ is defined by strict inequality constraints.
We explain the description of [26] using a mix of our notation and the notation in [26]. Let
$Q_{i}\coloneqq\left(\begin{smallmatrix}A_{i}&b_{i}\\ b_{i}^{\intercal}&c_{i}\end{smallmatrix}\right),\qquad\forall i=1,2.$
For any $\gamma\in{\mathbb{R}}^{2}$, define $Q[\gamma]\coloneqq\gamma_{1}Q_{1}+\gamma_{2}Q_{2}$. Let $\lambda_{i}(\cdot)$ denote the $i$th smallest eigenvalue of its argument. Then, under the assumption that $\hat{\cal S}$ is defined via strict inequalities, is nonempty, and $\operatorname{clconv}(\hat{\cal S})\neq{\mathbb{R}}^{n}$, [26] establishes that
$\operatorname{clconv}(\hat{\cal S})=\left\{x\in{\mathbb{R}}^{n}:\,[\gamma_{Y}^{(1)},q(x)]\leq 0,\,[\gamma_{Y}^{(2)},q(x)\leq 0]\right\},$
where $\gamma_{Y}^{(1)},\gamma_{Y}^{(2)}\in{\mathbb{R}}^{2}_{+}$ are vectors that generate the connected component (a cone) of
$\left\{\gamma\in{\mathbb{R}}^{2}_{+}:\,\lambda_{2}(Q[\gamma])\geq 0\right\}$
containing $\Gamma_{1}$. By the definition of $\gamma_{Y}^{(i)}$, the set $\left\{x\in{\mathbb{R}}^{n}:[\gamma_{Y}^{(i)},q(x)]\leq 0\right\}$ is either convex (and thus SOC representable in the homogenized space) or it is the union of an SOC and its negative (in the homogenized space) referred as a semi-cone. In the latter case, [26] further establishes that we can in fact restrict our attention to only one side of the associated nonconvex semi-cone to get a convex representation of $\operatorname{clconv}(\hat{\cal S})$.
To summarize, in order to test SDP exactness based on the characterization of [26], we will have to first compute $\operatorname{clconv}(\hat{\cal S})$ by computing $\gamma_{Y}^{(1)},\gamma_{Y}^{(2)}\in{\mathbb{R}}^{2}_{+}$ from $\left\{\gamma\in{\mathbb{R}}^{2}_{+}:\,\lambda_{2}(Q[\gamma])\geq 0\right\}$ and then check whether the set $\left\{x\in{\mathbb{R}}^{n}:\,[\gamma_{Y}^{(i)},q(x)]\leq 0,\,\forall i\in[2]\right\}$ is the same as the one given by $\left\{x\in{\mathbb{R}}^{n}:\,[\gamma^{(i)},q(x)]\leq 0,\,\forall i\in[2]\right\}$ obtained from $\Gamma_{1}$. In contrast to this our condition given in Lemma 4 simply requires us to compute $\Gamma_{1}=\operatorname{cone}\left\{\gamma^{(1)},\gamma^{(2)}\right\}$ and check for both $i\in[2]$ whether $\ker(A[\gamma^{(i)}])\cap b[\gamma^{(i)}]^{\perp}$ is nontrivial whenever $\gamma^{(i)}\in{\mathbb{R}}^{2}_{++}$; this can be done by solving a system of linear equations. In the rest of this remark we establish the equivalence of our condition and the one from [26] when SDP exactness holds.
To explain the presence of the second eigenvalue of $Q[\gamma]$, note that by assumption (of strict feasibility of $\hat{\cal S}$), there exists an $\hat{x}\in{\mathbb{R}}^{n}$ for which $q_{1}(\hat{x}),q_{2}(\hat{x})<0$ so that for any nonzero $\gamma\in{\mathbb{R}}^{2}_{+}$, we have
$\begin{pmatrix}\hat{x}\\ 1\end{pmatrix}^{\intercal}Q[\gamma]\begin{pmatrix}\hat{x}\\ 1\end{pmatrix}<0.$
We deduce that $Q[\gamma]$ has at least one negative eigenvalue for all nonzero $\gamma\in{\mathbb{R}}^{2}_{+}$. On the other hand, for any $\gamma\in\Gamma_{1}$, by Cauchy’s Eigenvalue Interlacing Theorem, we have that
$\lambda_{2}(Q[\gamma])\geq\lambda_{1}(A[\gamma])\geq 0$
so that $Q[\gamma]$ has exactly one negative eigenvalue for all nonzero $\gamma\in\Gamma_{1}$.
Now, outside of a few special edge cases (e.g., when $\operatorname{clconv}(\hat{\cal S})$ is defined by a single convex constraint), $\operatorname{clconv}(\hat{\cal S})=\hat{\cal S}_{\textup{SDP}}$ if and only if $\Gamma_{1}=\operatorname{cone}\left\{\gamma_{Y}^{(1)},\gamma_{Y}^{(2)}\right\}$, i.e., $\Gamma_{1}$ coincides with one of the connected components of $\left\{\gamma\in{\mathbb{R}}^{2}_{+}:\,\lambda_{2}(Q[\gamma])\geq 0\right\}$. The latter condition holds if and only if for both $i=1,2$, either $\gamma^{(i)}$ lies on the boundary of +${\mathbb{R}}^{2}_{+}$ or that $\lambda_{2}(Q[\gamma^{(i)}-\epsilon\hat{\gamma}])<0$ for all $\epsilon>0$ small enough, i.e., that the number of negative eigenvalues in $Q$ increases when moving out of $\Gamma_{1}$.
The remainder of this remark shows that this condition is equivalent to our requirement that $\ker(A[\gamma^{(i)}])\cap b[\gamma^{(i)}]^{\perp}$ is nontrivial if $\gamma^{(i)}\in{\mathbb{R}}^{2}_{++}$. To this end, suppose $\gamma^{(1)}\in{\mathbb{R}}^{2}_{++}$. The reasoning for $\gamma^{(2)}\in{\mathbb{R}}^{2}_{++}$ is analogous.
First, suppose that $\ker(A[\gamma^{(1)}])\cap b(\gamma^{(1)})^{\perp}$ is nontrivial and let $\xi$ be a unit vector from $\ker(A[\gamma^{(1)}])\cap b(\gamma^{(1)})^{\perp}$. Then, as $\gamma^{(1)}\in\Gamma_{1}$ is nonzero we deduce that $Q[\gamma^{(1)}]$ has exactly one negative eigenvalue. Furthermore,
$Q[\gamma^{(1)}]\begin{pmatrix}\xi\\ 0\end{pmatrix}=\begin{pmatrix}A[\gamma^{(1)}]\xi\\ b[\gamma^{(1)}]^{\intercal}\xi\end{pmatrix}=0$
so that $(\xi,0)$ is in $\ker(Q[\gamma^{(1)}])$ and hence $\lambda_{2}(Q[\gamma^{(1)}])=0$. Let $(\eta,t)$ denote a unit eigenvector corresponding to $\lambda_{1}(Q[\gamma^{(1)}])$. Our goal is to show that $\lambda_{2}(Q[\gamma^{(1)}-\epsilon\hat{\gamma}])<0$ for all $\epsilon>0$ small enough. By the variational characterization of $\lambda_{2}$, it suffices to show that the quadratic form induced by $Q[\gamma^{(1)}-\epsilon\hat{\gamma}]$ is negative for all $\epsilon>0$ small enough on some two-dimensional subspace (other than at zero). We will use the subspace spanned by $(\xi,0)$ and $(\eta,t)$ for this purpose. Let ${\mathbf{S}}^{1}\subseteq{\mathbb{R}}^{2}$ denote the unit circle and let $\epsilon>0$. Abusing notation, we will identify $v\in{\mathbf{S}}^{1}\subseteq{\mathbb{R}}^{2}$ with the unit vector $v_{1}(\xi;0)+v_{2}(\eta;t)$. Then,
$\lambda_{2}(Q[\gamma^{(1)}-\epsilon\hat{\gamma}])$ $\leq\max_{v\in{\mathbf{S}}^{1}}\left\{v^{\intercal}Q[\gamma^{(1)}-\epsilon\hat{\gamma}]v\right\}$
$=\max_{v\in{\mathbf{S}}^{1}}\left\{\begin{array}[]{l}\left(v_{1}^{2}\xi^{\intercal}A[\gamma^{(1)}-\epsilon\hat{\gamma}]\xi\right)\\ \quad+2v_{1}v_{2}\begin{pmatrix}\xi\\ 0\end{pmatrix}^{\intercal}Q[\gamma^{(1)}-\epsilon\hat{\gamma}]\begin{pmatrix}\eta\\ t\end{pmatrix}\\ \quad+v_{2}^{2}\begin{pmatrix}\eta\\ t\end{pmatrix}^{\intercal}Q[\gamma^{(1)}-\epsilon\hat{\gamma}]\begin{pmatrix}\eta\\ t\end{pmatrix}\end{array}\right\}$
$=\max_{v\in{\mathbf{S}}^{1}}\left\{\begin{array}[]{l}-\epsilon v_{1}^{2}\xi^{\intercal}A[\hat{\gamma}]\xi-2\epsilon v_{1}v_{2}\begin{pmatrix}\xi\\ 0\end{pmatrix}^{\intercal}Q[\hat{\gamma}]\begin{pmatrix}\eta\\ t\end{pmatrix}\\ \quad+v_{2}^{2}\lambda_{1}(Q[\gamma^{(1)}])-\epsilon v_{2}^{2}\begin{pmatrix}\eta\\ t\end{pmatrix}^{\intercal}Q[\hat{\gamma}]\begin{pmatrix}\eta\\ t\end{pmatrix}\end{array}\right\}$
$=\max_{v\in{\mathbf{S}}^{1}}v^{\intercal}\begin{pmatrix}-\epsilon\xi^{\intercal}A[\hat{\gamma}]\xi&\Theta(\epsilon)\\ \Theta(\epsilon)&\lambda_{1}(Q[\gamma^{(1)}])+\Theta(\epsilon)\end{pmatrix}v.$
Here, the $\Theta(\epsilon)$ denote functions of $\epsilon$ with magnitude bounded by $O(\epsilon)$. Thus, as $A[\hat{\gamma}]\succ 0$ and $\lambda_{1}(Q[\gamma^{(1)}])<0$, we deduce that the two-by-two matrix in the final line is negative definite so that $\lambda_{2}(Q[\gamma^{(1)}-\epsilon\hat{\gamma}])<0$ for all $\epsilon>0$ small enough.
In the reverse direction, suppose $\lambda_{2}(Q[\gamma^{(1)}-\epsilon\hat{\gamma}])<0$ for all $\epsilon>0$ small enough. By continuity of eigenvalues and the fact that $\lambda_{2}(Q[\gamma^{(1)}])\geq\lambda_{1}(A[\gamma^{(1)}])=0$, it must hold that $\lambda_{2}(Q[\gamma^{(1)}])=0$. We may then invoke the equality case of Cauchy’s Eigenvalue Interlacing Theorem [13, Theorem 4.3.17] to conclude that $\ker(A[\gamma^{(1)}])\cap b[\gamma^{(1)}]^{\perp}$ is nontrivial.